In a study at West Virginia University Hospital, researchers investigated smoking behavior of cancer patients to create a program to help patients stop smoking. They published the results in Smoking Behaviors Among Cancer Survivors (January 2018, Journal of Oncology Practice). In this study, the researchers sent a 22-item survey to 1484 cancer patients. They collected demographic information (age, sex, ethnicity, zip code, level of education), clinical and smoking history, and information about quitting smoking.

Of the 1484 patients who were mailed surveys, 297 patients responded. For various reasons, researchers used only 267 of the completed surveys. 42 out of 144 female cancer patients reported being past smokers, and 78 out of 123 male cancer patients reported being past smokers. Calculate the difference between the corresponding sample proportions ^p1−^p2 (female minus male). Round the answer to 4 decimal places.
Answer #1

\small \widehat{p}_1 = 42/144 = 0.2917

\small \widehat{p}_2 = 78/123 = 0.6341

Hence,

Difference = 0.2917 - 0.6341 = -0.3425